OpenStudy (anonymous):

i need to find the equation of a tangent line to a given curve at a given value of x

OpenStudy (anonymous):

y-y1 = m(x-x1) where (x1,y1) is a point on the curve. Then take the derivative of the curve, plug in x1 to the derivative, and that is your slope (or m)

OpenStudy (anonymous):

yes , this part i know. Not sure if im doing it right though. looking for help walking through it.

OpenStudy (anonymous):

What's the equation and x-value given?

OpenStudy (anonymous):

y=x(cubed) - 25x + 5; x = 5

OpenStudy (anonymous):

Okay, so f(5) = 5^3 - 25*5 + 5 = 5, so the point is (5,5)

OpenStudy (anonymous):

i think my equation (solution) is y = 50x - 245

OpenStudy (anonymous):

yes i got that too, so im feeling better

OpenStudy (anonymous):

Then taking the derivative of f(x), we get f'(x) to be 3x^2 - 25, and plugging in x=5 for that, we get 50.

OpenStudy (anonymous):

So the equation of the tangent line is y-5 = 50(x-5), or y = 50x - 245, so you are correct

OpenStudy (anonymous):

boom!

OpenStudy (anonymous):

thanks so much

OpenStudy (anonymous):

No problem buddy

OpenStudy (anonymous):

are you any good with rates of change?

OpenStudy (anonymous):

I'm good with any calculus; what's the question?

OpenStudy (anonymous):

D(p) = -3p(squared) + 7p + 7, where p is price, find D'(5)

OpenStudy (anonymous):

totally lost on this stuff

OpenStudy (anonymous):

D(p) = -3p^2 + 7p + 7, so the derivative would be D'(p) = -6p + 7. Plugging in p=5 would give us D'(5) = -23

OpenStudy (anonymous):

okay i see what you did, why do we just throw out the 7p ?

OpenStudy (anonymous):

I didn't, I took the derivative of 7p, which gave me 7. I got rid of the 7 because when you take the derivative of a constant (any number without a variable), then it becomes 0.

OpenStudy (anonymous):

oh ok i get it. i seem to remember going over that in class. our professor is way behind and he's flying through the book at a rate that we can't keep up with.

OpenStudy (anonymous):

you going to be on here awhile?

OpenStudy (anonymous):

Yeah I'll take any other questions you have.

OpenStudy (anonymous):

I can explain whatever you need me to explain.

OpenStudy (carson889):

A simple way of doing the derivative is f(x) =a*x^n , f'(x) = n*a*x^(n-1). This only holds for simple examples and cannot be used to derive something like (x^2 + 2)^2,notice the exponent outside the brackets, which requires the chain rule.

OpenStudy (anonymous):

i want to try to bang out more of this homework in a way that I actually learn it so let me try a few more and get back to you.

OpenStudy (carson889):

So for 7p^1, to the power of one is implied, using the formula, a = 7, n = 1, the derivative is 7*1*p^(1-0) = 7*1*p^0 = 7*1*1 = 7

OpenStudy (anonymous):

ya carson thats all giberrish to me. i wish looking at this stuff made sense but it just makes my eyes cross and everything goes blurry.

OpenStudy (anonymous):

i do remember going over that though and its coming back to me now. i just need to keep pressing on through these problems so that something will stick

OpenStudy (anonymous):

you guys are awesome. i'll be back in a few im sure. want to try a few more on my own.

OpenStudy (anonymous):

okay, using the product rule to find the derivative where \[f(x)=(x^{2}-3x+2)(2x ^{3}-x ^{2}+4)\], is the solution \[10x ^{4}-24x ^{3}+21x ^{2}+4x-12\]

OpenStudy (anonymous):

-24x^3 should be -28x^3, but other than that, everything else is correct.

OpenStudy (anonymous):

what did i not multiply correctly to acheive that answer? im doing that in my head mostly, like john nash.

OpenStudy (anonymous):

Probably multiplied wrong somewhere, with problems like this, it's really easy to miss something or add wrong.

OpenStudy (anonymous):

okay can you walk me through this, i dont have an example in my book for finding marginal demand. D(p) =(6p+250) / (8p+19), what is the marginal demand D'(p)